OpenStudy (anonymous):

You have three $1 bills, four $5 bills, and two $10 bills in your wallet. You select a bill at random. Without replacing the bill, you choose a second bill at random. Find P($10 then $1).

OpenStudy (oaktree):

It's 2/9 times 3/9 = 2/27.

OpenStudy (oaktree):

Or is addition, and is multiplication. We want $10 AND then $1, so we multiply the separate probabilities.

OpenStudy (oaktree):

Without replacing the bill, and you didn't add right. 3+2+4=9, not 16.

OpenStudy (anonymous):

Suppose you choose a marble from a bag containing 4 red marbles, 2 white marbles, and 3 blue marbles. You return the first marble to the bag and then choose again. Find P(red and blue). A. 4/3 B. 2/3 C. 7/9 D. 4/27

OpenStudy (oaktree):

Again, we multiply. 4/10 times 3/10 = 2/5 x 3/10 = 6/50 = 3/25.

OpenStudy (anonymous):

when you see "and" you should think "times"